OpenStudy (anonymous):

Log6 푥+7 −log64= log6(2푥+1)

OpenStudy (anonymous):

Log6 x+7 −log64= log6(2x+1)

OpenStudy (jdoe0001):

\(\bf recall\implies log_xa-log_xb=log_x\left(\frac{a}{b}\right)\) thus do that on the left side then remove the logs from both sides